# CM points on Shimura curves via QM-equivariant isogeny volcanoes

### v2

@@ 1. Introduction
The restriction of the study of torsion of elliptic curves over number fields to the case of complex multiplication (CM) has seen considerable recent progress. In particular, work of Clark and Clark–Saia [Cl23, CS23], continuing a program of research in this area from the perspective of CM points on modular curves (see, e.g., [CCS13, BC20, CGPS22]), approaches the study of the CM locus on the modular curves $X_{0}(M,N)_{/\mathbb{Q}}$ and $X_{1}(M,N)_{/\mathbb{Q}}$ via a study of CM components of isogeny graphs of elliptic curves over $\overline{\mathbb{Q}}$. For $K$ an imaginary quadratic field and $\Delta=f^{2}\Delta_{K}$ the discriminant of the $\mathfrak{o}(f)$ of conductor $f$ in $K$, let $j_{\Delta}\in X(1)_{/\mathbb{Q}}$ denote the closed point corresponding to elliptic curves with CM by the order of discriminant $\Delta$. This work results, for instance, in a description of all points in the fiber of the natural map $X_{0}(M,N)_{/\mathbb{Q}}\rightarrow X(1)_{/\mathbb{Q}}$ over $j_{\Delta}$. This description provides the list of residue fields of $\Delta$-CM points on the first curve, along with a count of closed points in this fiber with each specified residue field.
In this paper, we study the Shimura curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ and $X_{1}^{D}(N)_{/\mathbb{Q}}$ parametrizing abelian surfaces with quaternionic multiplication (QM) by the indefinite quaternion algebra $B$ over $\mathbb{Q}$ of discriminant $D$, along with certain specified level structure. Our main result allows for a similar description of the CM loci on these curves. In particular, for the first family we have a result of the following form. See §8.3 for details on the referenced algorithm.
@@
The $D=1$ case of this theorem results from the work of [Cl23] and [CS23]. For all $D$, we show the following: suppose that $x\in X_{0}^{D}(N)_{/\mathbb{Q}}$ has CM by the order $\mathfrak{o}(f)$ of conductor $f$ in the imaginary quadratic field $K$. The residue field $\mathbb{Q}(x)$ is then either a ring class field $K(f^{\prime})$ for some $f^{\prime}$ with $f\mid f^{\prime}\mid Nf$, or is isomorphic to an index $2$ subfield of such a field $K(f^{\prime})$. The ramification index mentioned in the algorithm is always $1$ when the CM order has discriminant $f^{2}\Delta_{K}=\Delta<-4$. In general, this index is at most $3$. This algorithm, outlined in Algorithm 8.2, has been implemented, and is publicly available at [Rep] along with [Magma] code for all other computations described in this paper.
@@
The ability to transition to information about the $\mathfrak{o}$-CM locus on $X_{1}^{D}(N)_{/\mathbb{Q}}$ is explained in §9, in which we prove the following result.
@@
Suppose that $x\in X_{0}^{D}(N)_{/\mathbb{Q}}$ is a point with CM by the imaginary quadratic order of discriminant $\Delta$. Let $\pi:X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ denote the natural morphism.
@@
    In all cases, the scheme-theoretic fiber of $\pi$ over $x$ consists of a single closed point.
@@
    If $\Delta<-4$, if $x$ is a ramified point of the map $X_{0}^{D}(N)_{/\mathbb{Q}}\rightarrow X^{D}(1)_{/\mathbb{Q}}$ or if $N\leq 3$, then $\pi$ is unramified over $x$.
@@
    If $\Delta<-4$, then the fiber of $\pi$ over $x$ consists of a single closed point.
  4. (4)
    If $N\geq 4$ and $x$ is unramified with respect to the natural map to $X^{D}(1)_{/\mathbb{Q}}$, then we have
    $e_{\pi}(x)=\begin{cases}2\quad&\textnormal{ if }\Delta=-4,\\ 3\quad&\textnormal{ if }\Delta=-3,\end{cases}\quad\textnormal{ and }\quad f_{\pi}(x)=\begin{cases}\phi(N)/4\quad&\textnormal{ if }\Delta=-4,\\ \phi(N)/6\quad&\textnormal{ if }\Delta=-3,\end{cases}$
    for the ramification index and residual degree of $x$, respectively, with respect to $\pi$.
Our work allows for a determination of all primitive residue fields and primitive degrees of $\mathfrak{o}$-CM points on $X_{0}^{D}(N)_{/\mathbb{Q}}$, as discussed in §8.4. Here, by a primitive residue field (respectively, a primitive degree) of an $\mathfrak{o}$-CM point on $X_{0}^{D}(N)_{/\mathbb{Q}}$ we mean one that does not properly contain (respectively, does not properly divide) that of another $\mathfrak{o}$-CM point on the same curve. An abridged version of the main result here is as follows, with Theorem 8.3 providing the complete result:
@@
Let $\mathfrak{o}(f)$ denote the order of conductor $f$ in an imaginary quadratic field $K$, and let $B$ denote the indefinite quaternion algebra over $\mathbb{Q}$ of discriminant $D$. Suppose that $K$ splits $B$, and that $N$ is a positive integer relatively prime to $D$, with prime-power factorization $N=\ell_{1}^{a_{1}}\cdots\ell_{r}^{a_{r}}$. Then one of the following occurs:
@@
    There are exactly $2$ primitive residue fields of such points, with one of the same form as $L$ in part (1) and the other being a ring class field of the form $K(Cf)$ with $C<H$.
@@
Let $\mathcal{F}$ be the explicit set of all $393$ pairs $(D,N)$ appearing in either Table 1 or in the file unknown_sporadics.m in [Rep]. If $(D,N)\not\in\mathcal{F}$ consists of a quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ which is relatively prime to $D$, then $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point. If $(D,N)$ is such a pair with
@@
The appearance of the pair $(91,5)$ in this result comes down to the fact that while $X_{0}^{91}(5)_{/\mathbb{Q}}$ has a sporadic CM point of degree $2$, the curve $X_{0}^{91}(5)_{/\mathbb{Q}}$ has $4$ as the least degree of a CM point. See Theorem 10.9 (4) for details.
Our work determining residue fields of CM points on $X_{0}^{D}(N)_{/\mathbb{Q}}$ can be viewed as a generalization of prior work on the Diophantine arithmetic of Shimura curves via an alternate approach (specifically work of Jordan [Jor81] and González–Rotger [GR06] – see Theorem 2.8). Of course, our results are also aimed towards better understanding the torsion of low-dimensional abelian varieties over number fields, via restriction to a case with extra structure. On this point, the question of which number fields admit abelian surfaces with certain specified rational torsion subgroups is closely related to our results, just as in the classical modular curve case. A result of Jordan (see Theorem 2.6) clarifies this relationship.
Unlike the modular curves $X_{0}(N)_{/\mathbb{Q}}$, the curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ for $D>1$ have no cusps. For this reason, understanding the CM points and, more generally, low degree points on Shimura curves may be of greater interest. Additionally, while our approach is in analogy to that of [Cl23] and [CS23] in the modular curve case, there are interesting deviations and difficulties arising in this work due to technical differences in the $D>1$ case. For example, fields of moduli are not necessarily fields of definition in the case of QM abelian surfaces, as noted in Jordan’s Theorem 2.6.
Further, while the field of moduli $\mathbb{Q}(x)$ of any CM point $x\in X(1)_{/\mathbb{Q}}$ has a real embedding, a result of Shimura [Sh75, Thm. 0] states that $X^{D}(1)_{/\mathbb{Q}}$ has no real points for $D>1$. This fact is inherent in fundamental differences between our general study and the $D=1$ case. It also opens the door for the potential of Hasse principal violations by Shimura curves, which has been a subject of significant study (see, e.g., [Cl09, CSt18, RSY05, SS16]). If one aims to study the Hasse principle for Shimura curves over some fixed number field (respectively, over a fixed degree), then studying the CM points rational over that field (respectively, over number fields of that degree) seems to be a natural initial point of investigation, and so our results may be of interest in that direction.
@@ Acknowledgments
We thank Pete L. Clark for initially suggesting this project, which served as the author’s main Ph.D. thesis work. We are very grateful for Clark’s excellent advising, and for innumerable helpful conversations. We also thank Dino Lorenzini, Oana Padurariu, Ciaran Schembri and John Voight for helpful comments and encouragement. Much gratitude goes as well to the anonymous referee for a careful reading of, and detailed feedback on, an earlier version of this work. This feedback led to many expository and mathematical improvements, including a strengthening of our results on sporadic points.
@@ 2.1. Shimura curves
on $\mathcal{O}$. We refer to $\mu$ as a principal polarization on $\mathcal{O}$
We start by defining the moduli spaces we are considering, and discussing the moduli interpretations of those families of particular interest to us in this study. Let $\mathcal{O}^{1}$ denote the units of reduced norm $1$ in $\mathcal{O}$, which we realize as embedded in $\text{SL}_{2}(\mathbb{R})$ via $\Psi$. The subgroup $\Gamma^{D}(1):=\Psi(\mathcal{O}^{1})\subset\text{SL}_{2}(\mathbb{R})$ is discrete, and is cocompact if and only if $D>1$. Via the action of this subgroup on the upper-half plane $\mathbb{H}$ we define over $\mathbb{C}$ the Shimura curve
$X^{D}(1)=X_{\Gamma^{D}(1)}:=\Gamma^{D}(1)\backslash\mathbb{H}.$
The $D=1$ case of $B\cong M_{2}(\mathbb{Q})$ recovers the modular curve setting $X^{1}(1)\cong Y(1)\cong\mathbb{A}^{1}_{\mathbb{C}}$. [^1] We do indeed mean $Y(1)$ rather than $X(1)$ here. Some may prefer the convention that we compactify by adding a cusp if $D=1$, which would yield $X^{1}(1)=X(1)$. The point here is that the classical modular curve case (compactified or not, up to convention) is recovered when $D=1$. The choice of convention will not affect anything we do in this paper, as we assume $D>1$ in the next sentence. Assume now that $D>1$, in which case $X^{D}(1)$ is a compact Riemann surface. For any $z\in\mathbb{H}$, we get a rank $4$ lattice $\Lambda_{z}$ via the action of $\mathcal{O}$ on $(1,z)\in\mathbb{C}^{2}$ via the embedding $\Psi$ above:
$\Lambda_{z}:=\mathcal{O}\cdot\begin{pmatrix}1\\ z\end{pmatrix}\subseteq\mathbb{C}^{2}.$
@@
For an abelian variety $A$ over a field $F$, by $\text{End}(A)$ we mean the ring of endomorphisms defined over a base field $F$. For an extension $F\subseteq L$, we will write $A_{L}$ for the base change of $A$ to $L$ and $\text{End}(A_{L})$ for the ring of endomorphisms rational over $L$.
More generally, if $\Gamma\leq\Gamma^{D}(1)\subseteq\text{SL}_{2}(\mathbb{R})$ is an arithmetic Fuchsian group, we can consider the curve $\ΓX_{\Gamma}=\Gamma\backslash\mathbb{H}$, and for $\Gamma^{\prime}\leq\Gamma$ there is a corresponding covering of curves $X_{\Gamma^{\prime}}\rightarrow X_{\Gamma}$. Our focus will be on the families of Shimura curves $X_{0}^{D}(N)$ and $X_{1}^{D}(N)$, for $N$ a positive integer with $\text{gcd}(D,N)=1$, with $X^{D}(1)=X^{D}_{0}(1)=X^{D}_{1}(1)$ being a special case of each.
@@
That is, $X_{0}^{D}(N)(\mathbb{C})=\Gamma^{D}_{0}(N)\backslash\mathcal{H}.$ Equivalently, fixing a level $N$ Eichler order $\mathcal{O}_{N}$ in $B$, the curve $X_{0}(N)$ can be described, in the manner mentioned above, as that associated to the arithmetic group of units of reduced norm $1$ in $\mathcal{O}_{N}$. The Shimura curve $X_{1}^{D}(N)$ corresponds to the compact, open subgroup
@@
For $N$ coprime to $D$, the notion of “level $N$-structure” is group-theoretically just as in the modular curve case. In particular, we have a natural modular covering map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ which is an isomorphism for $N\leq 2$, and has degree $\phi(N)/2$ for $N\geq 3$ where $\phi$ denotes the Euler totient function. We now recall moduli interpretations for these families of Shimura curves as provided in, for example, [Buz97, §3].
@@
The Shimura curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ is a coarse moduli space with the following equivalent moduli interpretations:
@@
These interpretations hold for any choice of principal polarization $\mu$ of $\mathcal{O}$. Moreover, if $\mu$ and $\mu^{\prime}$ are two such polarizations they both induce the same coarse moduli space $X_{0}^{D}(N)_{/\mathbb{Q}}$ up to isomorphism, as follows from the fact that their points over $\mathbb{C}$ are in bijection and thus they have the same canonical model over $\mathbb{Q}$ as obtained in [Sh67, Main Thm. 1]. (Of course, the moduli interpretation does depend on $\mu$, and we refer to [Rot04, Prop. 4.3] for more on how the corresponding spaces fit into the moduli space of principally polarized abelian surfaces.) Because a principal polarization $\lambda$ on a pair $(A,\iota)$ is canonically determined from a fixed $\mu$, moving forward we will suppress polarizations and refer simply to QM abelian surfaces $(A,\iota)$. By the same point, the condition on the polarizations in the definition of a QM-cyclic $N$-isogeny is redundant; it follows from the QM-equivariant condition.

### v3

@@ 1. Introduction
The restriction of the study of torsion of elliptic curves over number fields to the case of complex multiplication (CM) has seen considerable recent progress. In particular, work of Clark and Clark–Saia [Cl22, CS22], continuing a program of research in this area from the perspective of CM points on modular curves (see, e.g., [CCS13, BC20, CGPS22]), approaches the study of the CM locus on the modular curves $X_{0}(M,N)_{/\mathbb{Q}}$ and $X_{1}(M,N)_{/\mathbb{Q}}$ via a study of CM components of isogeny graphs of elliptic curves over $\overline{\mathbb{Q}}$. For $K$ an imaginary quadratic field and $\Delta=f^{2}\Delta_{K}$ the discriminant of the order $\mathfrak{o}(f)$ of conductor $f$ in $K$, let $j_{\Delta}\in X(1)_{/\mathbb{Q}}$ denote the closed point corresponding to elliptic curves with CM by the order of discriminant $\Delta$. The work of [Cl22, CS22] results, for instance, in a description of all points in the fiber of the natural map $X_{0}(M,N)_{/\mathbb{Q}}\rightarrow X(1)_{/\mathbb{Q}}$ over $j_{\Delta}$. This description provides the list of residue fields of $\Delta$-CM points on the first curve, along with a count of closed points in this fiber with each specified residue field.
In this paper, we study the Shimura curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ and $X_{1}^{D}(N)_{/\mathbb{Q}}$ parametrizing abelian surfaces with quaternionic multiplication (QM) by the indefinite quaternion algebra $B$ over $\mathbb{Q}$ of discriminant $D$, along with certain specified level structure. Our main result allows for a similar description of the CM loci on these curves.
In particular, we show that if $x\in X_{0}^{D}(N)_{/\mathbb{Q}}$ has CM by the order $\mathfrak{o}(f)$ of conductor $f$ in the imaginary quadratic field $K$, then the residue field $\mathbb{Q}(x)$ is either a ring class field $K(f^{\prime})$ for some $f^{\prime}$ with $f\mid f^{\prime}\mid Nf$, or is isomorphic to an index $2$ subfield of such a field $K(f^{\prime})$. The ramification index of $x$ with respect to the natural map from $X_{0}^{D}(N)$ to $X_{0}^{D}(1)$ is always $1$ when the CM order has discriminant $f^{2}\Delta_{K}=\Delta<-4$. In general, this index is at most $3$. The work of this paper culminates in a determination of the residue fields and ramification indices of all CM points on $X_{0}^{D}(N)$, and putting together the casework based on the quaternion discriminant, level and CM order gives a result of the following form.
@@
There exists an algorithm which, given as input
  • an indefinite quaternion discriminant $D$ over $\mathbb{Q}$,
  • a positive integer $N$ coprime to $D$ and
  • an imaginary quadratic discriminant $\Delta=f^{2}\Delta_{K}$,
@@
  • a boolean is_fixed,
  • a positive integer $f^{\prime}$ (necessarily with $f\mid f^{\prime}$),
  • an integer $e\in\{1,2,3\}$ and
  • a positive integer $c$
such that there exist exactly $c$ closed $\mathfrak{o}$-CM points $x$ on $X_{0}^{D}(N)_{/\mathbb{Q}}$ with the following properties:
  • the residue field of $x$ over $K$ is $K(x)\cong K(f^{\prime})$, the ring class field of conductor $f^{\prime}$ associated to $K$,
  • $\mathbb{Q}(x)\cong K(f^{\prime})$ if is_fixed is False,
  • $[K(f^{\prime}):\mathbb{Q}(x)]=2$ if is_fixed is True and
  • $x$ has ramification index $e$ with respect to the natural map to $X_{0}^{D}(1)_{/\mathbb{Q}}$.
This algorithm, outlined in Algorithm 8.2, has been implemented, and is publicly available at [Rep] along with [Magma] code for all other computations described in this paper.
@@
The ability to transition to information about the $\mathfrak{o}$-CM locus on $X_{1}^{D}(N)_{/\mathbb{Q}}$ is explained in §9, in which we prove the following result. While this does not determine the list of residue fields of CM points on $X_{1}^{D}(N)$ in the vein of Theorem 1.1, it allows us to count all CM points on $X_{1}^{D}(N)$ of specified degree and list their corresponding CM orders. Otherwise put, this is enough data to determine, for a fixed discriminant $D$ and degree $d$, all levels $N$ such that there exists a QM abelian surface $(A,\iota)$ and a torsion point $P\in A(\overline{\mathbb{Q}})$ of order $N$ such that the induced point $[A,\iota,P]\in X_{1}^{D}(N)$ has residue field of degree $d$.
@@
Suppose that $x\in X_{0}^{D}(N)_{/\mathbb{Q}}$ is a point with CM by the imaginary quadratic order of discriminant $\Delta$. Let $\pi_{1}:X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ and $\pi_{0}:X_{0}^{D}(N)_{/\mathbb{Q}}$ denote the natural morphisms. The following hold:
@@
    The scheme-theoretic fiber of $\pi_{1}$ over $x$ consists of a single closed point.
@@
    If any of the following hold:
    • $\Delta<-4$,
    • $x$ is ramified with respect to $\pi_{0}$ or
    • $N\leq 3$,
    then $\pi_{1}$ is unramified over $x$.
@@
    If $N\geq 4$ and $x$ is unramified with respect to $\pi_{0}$, then in the $\Delta\in\{-3,-4\}$ case we have
    $e_{\pi_{1}}(x)=\begin{cases}2\quad&\textnormal{ if }\Delta=-4,\\ 3\quad&\textnormal{ if }\Delta=-3,\end{cases}\quad\textnormal{ and }\quad f_{\pi_{1}}(x)=\begin{cases}\phi(N)/4\quad&\textnormal{ if }\Delta=-4,\\ \phi(N)/6\quad&\textnormal{ if }\Delta=-3,\end{cases}$
    for the ramification index and residue degree of $x$, respectively, with respect to $\pi_{1}$.
We define a primitive residue field (respectively, a primitive degree) of an $\mathfrak{o}$-CM point on $X_{0}^{D}(N)_{/\mathbb{Q}}$ to be one that does not properly contain (respectively, does not properly divide) that of another $\mathfrak{o}$-CM point on the same curve. Our work allows for a determination of all primitive residue fields and primitive degrees of $\mathfrak{o}$-CM points on $X_{0}^{D}(N)_{/\mathbb{Q}}$, as discussed in §8.4. An abridged version of our main result on primitive residue fields and degrees is as follows, with Theorem 8.3 providing the complete result:
@@
Suppose that $K$ splits $B$, let $f$ be a positive integer, and let $N$ be a positive integer relatively prime to $D$ with prime-power factorization $N=\ell_{1}^{a_{1}}\cdots\ell_{r}^{a_{r}}$. One of the following occurs:
@@
    There are exactly $2$ primitive residue fields of such points, with one of the same form as $L$ in part (1) and the other being a ring class field of the form $K(Cf)$ with $C<H$ and $C\mid N$.
@@
Let $\mathcal{F}$ be the set of all $393$ pairs $(D,N)$ appearing in Table 1 or Table LABEL:table:unknowns_table. If $(D,N)\not\in\mathcal{F}$ consists of a quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ which is relatively prime to $D$, then $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point. If $(D,N)$ is such a pair with
@@
The appearance of the pair $(91,5)$ in this result comes down to the fact that while $X_{0}^{91}(5)_{/\mathbb{Q}}$ has a sporadic CM point of degree $2$, the curve $X_{1}^{91}(5)_{/\mathbb{Q}}$ has $4$ as the least degree of a CM point. See Theorem 10.9 (4) for details.
Our work determining residue fields of CM points on $X_{0}^{D}(N)_{/\mathbb{Q}}$ can be viewed as a generalization of prior work on the Diophantine arithmetic of Shimura curves via an alternate approach (specifically work of Jordan [Jor81] and González–Rotger [GR06] – see Theorem 2.8). Of course, our results are aimed towards better understanding the torsion of low-dimensional abelian varieties over number fields, via restriction to a case with extra structure. On this point, the question of which number fields admit abelian surfaces with certain specified rational torsion subgroups is closely related to our results, just as in the classical modular curve case. A result of Jordan (see Theorem 2.6) clarifies this relationship.
Unlike the modular curves $X_{0}(N)_{/\mathbb{Q}}$, the curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ for $D>1$ have no cusps. For this reason, understanding the CM points on Shimura curves may be of even greater interest, as they provide the most accessible examples of low-degree points and could afford techniques (see, e.g., [BT07]) for computing models in the absence of techniques involving expansions around cusps.
Additionally, while our approach is in analogy to that of [Cl22] and [CS22] in the modular curve case, there are interesting deviations arising in this work due to technical differences in the $D>1$ case. Namely, while the field of moduli $\mathbb{Q}(x)$ of any CM point $x\in X(1)_{/\mathbb{Q}}$ has a real embedding, a result of Shimura [Sh75, Thm. 0] states that $X^{D}(1)_{/\mathbb{Q}}$ has no real points for $D>1$. This fact also opens the door for the potential of Hasse principle violations by Shimura curves, which has been a subject of significant study (see, e.g., [Cl09, CSt18, RSY05, SS16]). If one aims to study the Hasse principle for Shimura curves over some fixed number field (respectively, over a fixed degree), then studying the CM points rational over that field (respectively, over number fields of that degree) seems to be a natural initial point of investigation, and so our results may be of interest in that direction.
@@ Acknowledgments
We thank Pete L. Clark for initially suggesting this project, which served as the author’s main Ph.D. thesis work. We are very grateful for Clark’s excellent advising, and for innumerable helpful conversations. We also thank Dino Lorenzini, Oana Padurariu, Ciaran Schembri and John Voight for helpful comments and encouragement. Much gratitude goes as well to the anonymous referee for a careful reading of, and detailed feedback on, earlier versions of this work. This feedback led to many expository and mathematical improvements, including a strengthening of our results on sporadic points.
@@ 2.1. Shimura curves
on $\mathcal{O}$. We refer to $\mu$ as a principal polarization on $\mathcal{O}$.
We start by defining the moduli spaces we are considering and discussing the moduli interpretations of those families of particular interest to us in this study. Let $\mathcal{O}^{1}$ denote the units of reduced norm $1$ in $\mathcal{O}$, which we realize as embedded in $\text{SL}_{2}(\mathbb{R})$ via $\Psi$. The subgroup $\Gamma^{D}(1):=\Psi(\mathcal{O}^{1})\subset\text{SL}_{2}(\mathbb{R})$ is discrete, and it is cocompact if and only if $D>1$. Via the action of this subgroup on the upper-half plane $\mathbb{H}$ we define over $\mathbb{C}$ the Shimura curve
$X^{D}(1):=\Gamma^{D}(1)\backslash\mathbb{H}.$
For $D=1$ we have $B\cong M_{2}(\mathbb{Q})$, which recovers the familiar modular curve setting. We are interested in the $D>1$ case, and so moving forward we make this assumption on $D$. This impies that $X^{D}(1)$ is a compact Riemann surface. For any $z\in\mathbb{H}$, we get a rank $4$ lattice $\Lambda_{z}$ via the action of $\mathcal{O}$ on $(z,1)\in\mathbb{C}^{2}$ via the embedding $\Psi$ above:
$\Lambda_{z}:=\mathcal{O}\cdot\begin{pmatrix}z\\ 1\end{pmatrix}\subseteq\mathbb{C}^{2}.$
@@
For an abelian variety $A$ over a field $F$, by $\text{End}(A)$ we mean the ring of endomorphisms defined over $F$. For an extension $F\subseteq L$, we will write $A_{L}$ for the base change of $A$ to $L$ and $\text{End}(A_{L})$ for the ring of endomorphisms rational over $L$.
More generally, if $\Gamma\leq\Gamma^{D}(1)\subseteq\text{SL}_{2}(\mathbb{R})$ is an arithmetic Fuchsian group, we can consider the curve $\Γ\Gamma\backslash\mathbb{H}$, and for $\Gamma^{\prime}\leq\Gamma$ there is a corresponding covering of curves $\Γ\Gamma^{\prime}\backslash\mathbb{H}\rightarrow\Gamma\backslash\mathbb{H}$. Our focus will be on the families of Shimura curves $X_{0}^{D}(N)$ and $X_{1}^{D}(N)$, for $N$ a positive integer with $\text{gcd}(D,N)=1$, with $X^{D}(1)=X^{D}_{0}(1)=X^{D}_{1}(1)$ being a special case of each.
@@
That is, $X_{0}^{D}(N)(\mathbb{C})=\Gamma^{D}_{0}(N)\backslash\mathbb{H}.$ Equivalently, fixing a level $N$ Eichler order $\mathcal{O}_{N}$ in $B$, the curve $X_{0}(N)$ can be described, in the manner mentioned above, as that associated to the arithmetic group of units of reduced norm $1$ in $\mathcal{O}_{N}$. The Shimura curve $X_{1}^{D}(N)$ corresponds to the compact, open subgroup
@@
Because we are assuming that $N$ is relatively prime to $D$, the notion of “level $N$-structure” is group-theoretically just as in the modular curve case. In particular, the natural modular map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ is a $(\mathbb{Z}/N\mathbb{Z})^{\times}/\{\pm 1\}$-cover. Hence, it is an isomorphism for $N\leq 2$ and it has degree $\phi(N)/2$ for $N\geq 3$, where $\phi$ denotes the Euler totient function. We now recall moduli interpretations for these families of Shimura curves as provided in, for example, [Buz97, §3].
@@
The Shimura curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ is isomorphic to the coarse moduli scheme associated to any of the following moduli problems:
@@
These interpretations hold for any choice of principal polarization $\mu$ of $\mathcal{O}$. That is, if $\mu$ and $\mu^{\prime}$ are two such polarizations then they both induce the same coarse moduli scheme $X_{0}^{D}(N)_{/\mathbb{Q}}$ up to isomorphism (as discussed, for example, in [Rot04, §6]). Of course, the exact moduli interpretation does depend on $\mu$, and we refer to [Rot04, Prop. 4.3] for more on how the corresponding spaces fit into the moduli space of principally polarized abelian surfaces. Because a principal polarization $\lambda$ on a pair $(A,\iota)$ is canonically determined from a fixed $\mu$, moving forward we will suppress polarizations and refer simply to QM abelian surfaces $(A,\iota)$. By the same point, the condition on the polarizations in the definition of a QM-cyclic $N$-isogeny is redundant; it follows from the QM-equivariant condition.